OpenStudy (anonymous):

Will give a medal! Donna teaches dance in groups of five students. She rents a hall for the dance classes at $450 per month. She charges $40 per hour per student. Donna teaches four groups of students, and her net earnings per month are $4,350. Part A: If each student learns dance for the same number of hours, write an equation that can be used to calculate the number of hours each student learns dance in a month. Part B: How many hours did each student learn dance each month? Show your work and justify each step.

OpenStudy (anonymous):

Part A: First you need to find the amount of money Donna earns without the deduction for the rent. 4,350+450=4800. Next you need to find how many students Donna has. She has four groups of five students, this means she has 20 overall students.

OpenStudy (anonymous):

@dtan5457 @jhannybean @nnesha @satellite73

OpenStudy (dtan5457):

kind of tricky wording but it's pretty simple. the total is $4800 so...

OpenStudy (dtan5457):

your going to divide by by $40 as it is per student to get

OpenStudy (anonymous):

120

OpenStudy (dtan5457):

120 hours , which must be divided by 5 to get 24 hours per student

OpenStudy (dtan5457):

part a:24 hrs

OpenStudy (anonymous):

but there are four groups of five so that makes 20 students?

OpenStudy (dtan5457):

it says groups of 5

OpenStudy (anonymous):

but there are four groups with five students in each...

OpenStudy (anonymous):

still here? @dtan5457

OpenStudy (dtan5457):

yea. i think your right. i had to read that over again. case of 20 students. 120/20=6 hours per student

OpenStudy (anonymous):

ok so what would the equation be?

OpenStudy (dtan5457):

$4800<base price $4800/40=20x

OpenStudy (dtan5457):

120=20x x=6

OpenStudy (anonymous):

is there a way to put that all in one equation?

OpenStudy (dtan5457):

$4800/$40=20x

OpenStudy (dtan5457):

i just solved it

OpenStudy (anonymous):

could you help me with another @dtan5457

OpenStudy (dtan5457):

ok tag me in the new question